# The JARVIS Infrastructure is All You Need for Materials Design

### v1

@@ 3.2.3 Transformer Models
In addition to GNNs, JARVIS explores the application of transformer-based models in materials science. Originally developed for natural language processing( NLP such as in ChemNLP [52]), Transformers have been adapted to handle the complex patterns found in material structures for both forward and inverse materials design such as in AtomGPT [38] and DiffractGPT [53] frameworks. Recent research has introduced models that generate atomic embeddings using Transformer architectures, leading to enhanced predictions of material properties. These models capture long-range dependencies and intricate relationships within the material data, offering a powerful tool for materials informatics.
By integrating these machine learning methodologies-fingerprinting techniques, Graph Neural Networks, and Transformer-based models—JARVIS provides a robust framework for the accelerated discovery and design of new materials.
3.3 Classical Methods and Force Fields
The JARVIS infrastructure integrates classical simulation methods and force fields to complement its quantum mechanical approaches, thereby providing a comprehensive toolkit for materials modeling.
@@ 3.3.1 Conventional Force Fields
A fundamental component of JARVIS is Classical Molecular Dynamics (MD), which utilizes empirical interatomic potentials to simulate the behavior of atoms and molecules over time. This method is particularly effective for studying the thermodynamic and kinetic properties of materials at finite temperatures and for modeling large systems that are computationally prohibitive for quantum mechanical methods. JARVIS-FF [54, 55] maintains a repository of classical force fields tailored for various materials, facilitating accurate MD simulations within its framework with Large-scale Atomic/Molecular Massively Parallel Simulator, General Utility Lattice Program (GULP), Assisted Model Building with Energy Refinement (AMBER) and similar MD software packages.
@@ 4.3 Carbon Capture and Sustainable Materials
Addressing environmental challenges, JARVIS integrates data on materials designed for carbon capture and sustainability. This includes information on metal-organic frameworks (MOFs), zeolytes, polymeric membranes and other porous materials that can adsorb carbon dioxide, contributing to efforts in mitigating climate change.
@@ 4.6 Defects and Interfaces
Understanding defects and interfaces/heterostructures is crucial for tailoring material properties for specific applications. JARVIS includes data on various defect structures and their properties such as vacancies, enabling researchers to predict how these imperfections can influence material behavior. This is particularly important in semiconductors, where defects can significantly impact electronic properties. Similarly, JARVIS provides databases, tools, webapps, and benchmarks on solid-molecule interface (such as for catalysts) and solid-solid interface (such as for heterostructures used in microelectronics) are crucial for materials design.
@@ 5.2 Interactive Web Applications
To enhance user engagement and accessibility, JARVIS offers a suite of web applications that provide intuitive interfaces for data retrieval and analysis. These tools enable researchers to explore datasets, perform simulations, and visualize results without the need for extensive computational resources, thereby lowering the barrier to entry for materials research.
@@ 5.4 Benchmarking and Validation
JARVIS has developed the JARVIS-Leaderboard, an open-source, community-driven platform designed to facilitate benchmarking across various materials design methods. This initiative allows for the comparison of computational predictions with experimental datasets, providing a framework for the performance evaluation of different computational approaches. By systematically validating methods against empirical data, JARVIS ensures the reliability and accuracy of its predictive models.
Through these initiatives, JARVIS not only advances materials science but also sets a standard for reproducibility and transparency in computational research.
@@ 6.1 Global User Base and Downloads
Since its inception in January 2017, JARVIS has garnered a substantial international user base, with over 150,000 users accessing its resources. The papers in JARVIS have been cited over 4000 times. The platform’s extensive databases, encompassing more than 80,000 materials and over a million calculated properties, have been downloaded over a million times, underscoring its widespread adoption and utility in the research community. Examples of impact in terms of paper citations, software and data downloads, and global user base are shown in Fig. 4.
@@ 9 Data Availability
The JARVIS data is available at https://jarvis.nist.gov/.
@@ 10 Acknowledgments
K.C thanks the National Institute of Standards and Technology for funding, computational, and data-management resources. This work was performed with funding from the CHIPS Metrology Program, part of CHIPS for America, National Institute of Standards and Technology, U.S. Department of Commerce. Certain commercial equipment, instruments, software, or materials are identified in this paper in order to specify the experimental procedure adequately. Such identifications are not intended to imply recommendation or endorsement by NIST, nor it is intended to imply that the materials or equipment identified are necessarily the best available for the purpose. This is an Official contribution of the National Institute of Standards and Technology; not subject to copyright in the United States.

### v2

@@ 3.2.3 Transformer Models
In addition to GNNs, JARVIS explores the application of transformer-based models in materials science. Originally developed for natural language processing( NLP such as in ChemNLP [54]), Transformers have been adapted to handle the complex patterns found in material structures for both forward and inverse materials design such as in AtomGPT [40], DiffractGPT [55] and MicroscopyGPT [56] frameworks. Recent research has introduced models that generate atomic embeddings using Transformer architectures, leading to enhanced predictions of material properties. These models capture long-range dependencies and intricate relationships within the material data, offering a powerful tool for materials informatics.
By integrating these machine learning methodologies-fingerprinting techniques, Graph Neural Networks, and Transformer-based models-JARVIS provides a robust framework for the accelerated discovery and design of new materials.
@@ 3.2.4 Conventional Force Fields
A fundamental component of JARVIS is Classical Molecular Dynamics (MD), which utilizes empirical interatomic potentials to simulate the behavior of atoms and molecules over time. This method is particularly effective for studying the thermodynamic and kinetic properties of materials at finite temperatures and for modeling large systems that are computationally prohibitive for quantum mechanical methods.
JARVIS-FF [57, 58] maintains a curated repository of classical force fields tailored for various materials, enabling accurate MD simulations with software packages such as the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) [17], General Utility Lattice Program (GULP) [59], and AMBER [60]. In its early development, JARVIS-FF built upon foundational work by Chandler Becker [61] and leveraged contributions from related repositories such as OpenKIM [62].
It is often difficult to determine a priori whether a given force field is suitable for a particular material or property, especially when ground truth data are unavailable or inconsistent. JARVIS-FF addresses this challenge by linking high-throughput FF calculations directly to high-fidelity DFT reference data in the JARVIS-DFT database, enabling 1:1 benchmarking across properties such as lattice constants, formation energies, phonon spectra, and elastic moduli. This benchmarking-oriented approach goes beyond merely hosting force fields by providing automated workflows (e.g., SLURM-compatible job scripts), ready-to-use LAMMPS input templates, and comprehensive metadata tracking. All simulation scripts and examples, including Google Colab notebooks, are openly accessible at https://github.com/JARVIS-Materials-Design/jarvis-tools-notebooks. This infrastructure empowers users to perform reproducible, scalable, and validated MD simulations tailored to their material systems of interest.
@@ 4.3 Carbon Capture and Sustainable Materials
Addressing environmental challenges, JARVIS integrates data on materials designed for carbon capture and sustainability. This includes information on metal-organic frameworks (MOFs), zeolites, polymeric membranes and other porous materials that can adsorb carbon dioxide, contributing to efforts in mitigating climate change.
@@ 4.6 Defects and Interfaces
Understanding defects and interfaces/heterostructures is crucial for tailoring material properties for specific applications [81, 82, 83]. JARVIS includes data on various defect structures and their properties such as vacancies, enabling researchers to predict how these imperfections can influence material behavior. This is particularly important in semiconductors, where defects can significantly impact electronic properties. Similarly, JARVIS provides databases, tools, webapps, and benchmarks on solid-molecule interface (such as for catalysts) and solid-solid interface (such as for heterostructures used in microelectronics) are crucial for materials design.
@@ 5.2 Interactive Web Applications
To enhance user engagement and accessibility, JARVIS offers a suite of web applications that provide intuitive interfaces for data retrieval and analysis. These tools enable researchers to explore datasets, perform simulations, and visualize results without the need for extensive computational resources, thereby lowering the barrier to entry for materials research. A list of such apps and brief description are given in Table 3.
Name URL Description
1.ALIGNN Property Predictor https://jarvis.nist.gov/jalignn/ Atomistic Line Graph Neural Network (ALIGNN) for rapid property prediction.
2.ALIGNN Force‑Field https://jarvis.nist.gov/jalignnff/ GNN‑based interatomic potential for fast structure optimisation.
3.Solar Cells https://jarvis.nist.gov/jarvissolar Estimates theoretical photovoltaic performance of a material.
4.Direct Air Capture https://jarvis.nist.gov/jdac Predicts $\mathrm{CO_{2}}$ adsorption isotherms for metal‑organic frameworks.
5.Scanning Tunnelling Microscopy https://jarvis.nist.gov/jarvisstm Generates Tersoff-Hamann STM images from crystal structures.
6.Scanning Transmission Electron Microscopy https://jarvis.nist.gov/jstem Simulates STEM images using the convolution approximation.
7.Heterostructure Builder https://jarvis.nist.gov/jarvish Creates interfaces/heterostructures with the Zur lattice‑matching algorithm.
8.Catalysis (Adsorption) https://jarvis.nist.gov/jcatalysis Predicts adsorption energies of molecules on catalytic substrates.
9.Visualization https://jarvis.nist.gov/jarvisviz Lightweight, in‑browser atomic structure visualiser.
10.JARVIS‑XRD https://jarvis.nist.gov/jxrd Computes theoretical X‑ray diffraction patterns.
11.JARVIS‑WTBH https://jarvis.nist.gov/jarviswtb Calculates properties from Wannier tight‑binding Hamiltonians.
12.JARVIS‑Battery https://jarvis.nist.gov/jbattery Predicts capacities and voltage profiles of battery electrodes.
13.JARVIS‑ML / CFID https://jarvis.nist.gov/jarvisml Property prediction with Classical Force‑Field Inspired Descriptors.
Table 3: Selected JARVIS web applications and their functionalities in JARVIS.
@@ 5.4 Benchmarking and Validation
JARVIS has developed the JARVIS-Leaderboard, an open-source, community-driven platform designed to facilitate benchmarking across various materials design methods as mentioned above. This initiative allows for the comparison of computational predictions with experimental datasets, providing a framework for the performance evaluation of different computational approaches. By systematically validating methods against empirical data, JARVIS ensures the reliability and accuracy of its predictive models. Through these initiatives, JARVIS not only advances materials science but also sets a standard for reproducibility and transparency in computational research.
@@ 6.1 Global User Base and Downloads
Since its inception in January 2017, JARVIS has garnered a substantial international user base, with over 150,000 users accessing its resources. The papers in JARVIS have been cited over 4000 times. The platform’s extensive databases, encompassing more than 90,000 materials and over a million calculated properties, have been downloaded over a million times, underscoring its widespread adoption and utility in the research community. Examples of impact in terms of paper citations, software and data downloads, and global user base are shown in Fig. 4.
@@ 9 Data Availability
The JARVIS data is available at websites https://jarvis.nist.gov/ and https://atomgpt.org/.
@@ 10 Acknowledgments
K.C. thanks the National Institute of Standards and Technology as well as the Johns Hopkins University for funding, computational, and data-management resources. This work was performed with funding from the CHIPS Metrology Program, part of CHIPS for America, National Institute of Standards and Technology, U.S. Department of Commerce. Certain commercial equipment, instruments, software, or materials are identified in this paper in order to specify the experimental procedure adequately. Such identifications are not intended to imply recommendation or endorsement by NIST, nor it is intended to imply that the materials or equipment identified are necessarily the best available for the purpose.